For each problem, find the values of c that satisfy Rolles Theorem. 8) y = sin (2x); (-1,7) A) (0) D) 31 4440
Answer #1

Qe Y = Sin (2x) . [-tan] Here y(-1) = 0 , y(«) 20 = 2.CH(2x). y = sin(2x Rolls theorem - tf (a) = f(b) then f(c) zo akc<b